OpenStudy (anonymous):

How do I find the x and y intercepts for y=4(x-2)squared

OpenStudy (anonymous):

set x = 0 y = 4 * (0-2) = -8 set y = 0 0 = 4(x-2)^2 0 = (x-2)^2 x = 2

OpenStudy (anonymous):

when solving x what happened to the 4?

OpenStudy (anonymous):

you divide both sides by 4 and 0 divided by 4 is 0

OpenStudy (anonymous):

Also in finding y what happened to the square? y=4(x-2)^2

OpenStudy (anonymous):

nothing, it means that the curve just touches the x axis instead of intersecting it twice. draw it out, there are online autographs available and you'll know what i mean.

OpenStudy (anonymous):

I worked it out as below. Where have I gone wrong y=4*(0-2)^2 y=4*(0-2)(0-2) y=-8 -2 y=16

OpenStudy (anonymous):

you are right, i forgot about the squared bit

OpenStudy (anonymous):

Thanks heaps for your help